OpenStudy (anonymous):

The magnetic induction at a distance of 0.1 m from a straight wire through which 10A current flows is: A) 0.2 x 10^‐5T B) 2 x 10^ ‐5T           C) 0.02 x 10^‐5T D) 0.002 x 10^‐5T

OpenStudy (anonymous):

for a straight wire magentic field B at a distance of 0.1 m away is given by, B = mu(I)/2(pi)r where I is the current and r the radial distance from the wire.plugging the values in the equation gives you B=10^(-7) *2*10/0.1 (as mu/2(pi) = 2*10^(-7)) B=2*10^-5T